OpenStudy (anonymous):

The graph is of x/x^2-x has the discontinuities at: Non-removable at x = 0 Removable at x = 0 Removable at x = 0; Non-removable at x =1 Non-removable at x = 0; Removable at x = 1 The graph is continuous for all real numbers.

OpenStudy (anonymous):

Removable at x = 0; Non-removable at x =1 Do you need explanation?

OpenStudy (anonymous):

Are you positive? @eliassaab

OpenStudy (anonymous):

\[ \frac x{x(x-1)}=\frac 1{x-1} \]

OpenStudy (anonymous):

Yes, I am positive.